Three urns U portions: 1,2 and 3 contain red and and yellow balls in the following pro- Red43 4 Yellow 23 7 If a ball is withdrawn from each urn, what is the probability that they are all of th<e same color?
Answer #1

Required probability = P(all balls drawn are of the same color)

= P(all drawn balls are Red) + P(all drawn balls are Yellow)

= 11 + 6 × 6 × її = 22 0.2273